OpenStudy (samsan9):

solve this polynomial equation. 15x^2-119x^2-10x+16=0

OpenStudy (imstuck):

is that 2 x^2 terms in there?

OpenStudy (imstuck):

\[15x ^{2}-119x ^{2}-10x+16\]is what you have typed

OpenStudy (anonymous):

15x^3 I think

OpenStudy (imstuck):

me too

OpenStudy (samsan9):

yes and with 15^3

OpenStudy (samsan9):

sorry i wrote it wrong

OpenStudy (samsan9):

\[15x^3-119x^2-10x+16=0\]

OpenStudy (imstuck):

(x-8) is the first root. Now let's find the rest!

OpenStudy (imstuck):

once we find the first factor of (x-8) we have the following polynomial left over to factor:\[15x ^{2}+x-2\]I would do this using the quadratic formula cuz it looks hairy.

OpenStudy (samsan9):

so you have to factor it out first?

OpenStudy (imstuck):

The other factors are this:

OpenStudy (imstuck):

\[(x-\frac{ 1 }{ 3 })(x+\frac{ 2 }{ 5 })\]giving you a whole factorization of:\[(x-8)(x-\frac{ 1 }{ 3 })(x+\frac{ 2 }{ 5 })\]

OpenStudy (imstuck):

factoring is the only way to solve these.

OpenStudy (imstuck):

I used synthetic division to find the first factor by the rules of the Rational Roots Theorem.

OpenStudy (samsan9):

ok and then what?

OpenStudy (imstuck):

watch. i will do it with 8 so you can see, ok?

OpenStudy (imstuck):

That 0 remainder tells us that (x-8) is a root.